OpenStudy (anonymous):

Multiply (s+1/2)(s+1/5)

OpenStudy (anonymous):

I got (2s+1)(5s+1)/10

OpenStudy (anonymous):

s^2+7s/10+1/10

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

is it only 1/2 and 1/5 or the whole expression being divided by 2 and 5

OpenStudy (anonymous):

s+1/2 is in one () and s+1/5 is in the other ().

OpenStudy (anonymous):

I mean is it [(s+1)/2][(s+1)/5

OpenStudy (anonymous):

If I were to multiply this expression I would FOIL it, as in first, outer,inner, and last variables. So (s+1/2)(s+1/5)= you would multiply the first variables of each equation. s*s=s^2 then you would multiply the outer variables of each equation s*1/5=1/5s Then the inner variables s*1/2=1/2s and finally the last variables 1/2*1/5=1/10 So your finished equation is s^2+1/5s+1/2s+1/10 which can be simplified further into (s^2)+(7/10s)+(1/10)

OpenStudy (anonymous):

\[\left(s+\frac{1}{5}\right) \left(s+\frac{1}{2}\right)=s^2+\frac{7 s}{10}+\frac{1}{10}=\frac{1}{10} \left(10 s^2+7 s+1\right) \]